1.9.given A "\begin{bmatrix}
 1 & 1& -1 \\
 2 & 2 & 3 \\
 4 & 0 &0 \\
 
\end{bmatrix}" B="\begin{bmatrix}
 1 & 1 & -1 \\
 2 & 2 & 3 \\
 4 & 0 & 0 \\

\end{bmatrix}"


1.9.1. find -A -1 +3BT

1.9.2. Find B-1+(AT+A-1)



1
Expert's answer
2022-04-05T16:23:29-0400

Remind that for "3\times3" matrix "T=\left(\begin{array}{lll}a_{11}&a_{12}&a_{13}\\a_{21}&a_{22}&a_{23}\\a_{31}&a_{32}&a_{33}\end{array}\right)" the inverse can be computed using the formula: "T^{-1}=\frac{1}{|T|}Adj\,\,T", where "|T|" is the determinant of matrix "T". It has the form: "|T|=a_{11}a_{22}a_{33}+a_{21}a_{13}a_{32}+a_{31}a_{12}a_{23}-a_{11}a_{23}a_{32}-a_{12}a_{21}a_{33}-a_{31}a_{13}a_{22}" ; "Adj\,\,T=\left(\begin{array}{lll}A_{11}&A_{12}&A_{13}\\A_{21}&A_{22}&A_{23}\\A_{31}&A_{32}&A_{33}\end{array}\right)" and "A_{11}=\left|\begin{array}{lll}a_{22}&a_{23}\\a_{32}&a_{33}\end{array}\right|", "A_{12}=-\left|\begin{array}{lll}a_{12}&a_{13}\\a_{32}&a_{33}\end{array}\right|", "A_{13}=\left|\begin{array}{lll}a_{12}&a_{13}\\a_{22}&a_{23}\end{array}\right|", "A_{21}=-\left|\begin{array}{lll}a_{21}&a_{23}\\a_{31}&a_{33}\end{array}\right|", "A_{22}=\left|\begin{array}{lll}a_{11}&a_{13}\\a_{31}&a_{33}\end{array}\right|", "A_{23}=-\left|\begin{array}{lll}a_{11}&a_{13}\\a_{21}&a_{23}\end{array}\right|" , "A_{31}=\left|\begin{array}{lll}a_{21}&a_{22}\\a_{31}&a_{32}\end{array}\right|", "A_{32}=-\left|\begin{array}{lll}a_{11}&a_{12}\\a_{31}&a_{32}\end{array}\right|", "A_{33}=\left|\begin{array}{lll}a_{11}&a_{12}\\a_{21}&a_{22}\end{array}\right|."

For the given matrices we have:

a). "A^{-1}=\left(\begin{array}{lll}0&0&\frac{1}{4}\\\frac{3}{5}&\frac{1}{5}&-\frac{1}{4}\\-\frac{2}{5}&\frac{1}{5}&0\end{array}\right)"; "B^{\top}=\left(\begin{array}{lll}1&2&4\\1&2&0\\-1&3&0\end{array}\right)";

"-A^{-1}+3B^{\top}=\left(\begin{array}{lll}3&6&\frac{47}{4}\\\frac{12}{5}&\frac{29}{5}&\frac{1}{4}\\-\frac{13}{5}&\frac{44}{5}&0\end{array}\right)";

b). "B^{-1}=A^{-1}" . We receive:

"B^{-1}+(A^{\top}+A^{-1})=2A^{-1}+A^{\top}=\left(\begin{array}{lll}1&2&\frac{9}{2}\\\frac{11}{5}&\frac{12}{5}&-\frac{1}{2}\\-\frac{9}{5}&\frac{17}{5}&0\end{array}\right)".